3. Is it possible to design a rectangular mango grove whose length is twice its breadth, and the area is 800 m2? If so, find its length and breadth.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Let the breadth of mango grove be l. Length of mango grove will be 2l. Area of mango grove = (2l) (l)= 2l2 2l2 = 800 l2 = 800/2 = 400 l2 – 400 =0 Comparing the given equation with ax2 + bx + c = 0, we get a = 1, b = 0, c = 400 As we know, Discriminant = b2 – 4ac => (0)2 – 4 × (1) × ( – 400) = 1600 Here, b2 – 4ac > 0 Thus, the equation will have real roots. And hence, the desired rectangular mango grove can be designed. l = ±20 As we know, the value of length cannot be negative. Therefore, breadth of mango grove = 20 m Length of mango grove = 2 × 20 = 40 m